# Discrete Brunn–Minkowski inequality for subsets of the cube

### v1

@@ Abstract.
We show that for all subsets $A_{1},A_{2},\dots,A_{n}\subseteq\{0,1\}^{d}$, we have
$|A_{1}+A_{2}+\dots+A_{n}|\geqslant\prod_{i=1}^{n}|A_{i}|^{(\log_{2}(n+1))/n}.$
In addition, we show that for all $A,B\subseteq\{0,1,2\}^{d}$ we have
@@
holds for all $p\in(1,\infty)$, where $q=\frac{p}{p-1}$ is the conjugate exponent of $p$. All three estimates are dimension free with the best possible exponents. We also discuss applications to various related problems.
@@ 2. Choosing $A_{1}=\ldots=A_{n}=\mathcal{C}_{m}^{d}$ we obtain
3. The main result in [17] states that $p_{2,1}=\frac{\log(3)}{\log(4)}$, i.e., we have equality in (1.5) in the case $(n,m)=(2,1)$.
@@ 4. For $m\geqslant 3$ let $\tau_{m}$ be the solution of the equation
In the next theorem we show that we have equality in (1.5) for the pairs $(n,m)=(n,1)$ for all $n\geqslant 1$.
Theorem 1.1.
[…]
In particular, for all subsets $A_{j}\subseteq\{0,1\}^{d}$, we have
with the best possible power $p=\frac{\log_{2}(n+1)}{n}$.
The inequality (1.7) follows from (1.6) by taking $f_{j}=1_{A_{j}}$ for each $j=1,\ldots,n$. A related inequality to (1.6) was established by Ivanisvili [8, Theorem 1].
Next we affirmatively answer the question asked in [11] which also confirms the case $m=2$ of the question in [1].
@@
[…]
@@
Our Theorem 1.5 can be seen as an extension of a result obtained by Green and Tao [6, Proposition 4.1], for $k=2$.
@@
The additive energy $E(A)$ of a finite subset $A$ of an additive group $G$ is defined as $E(A):=|\{(a_{1},a_{2},a_{3},a_{4})\in A^{4};\ a_{1}+a_{2}=a_{3}+a_{4}\}|,$ (see [16]). It was proved by Kane and Tao [10, Theorem 7] that $E(A)\leqslant|A|^{\log_{2}6}$ for all $A\subseteq\{0,1\}^{d}$. For any $k\geqslant 2$, the $k-$additive energy $E_{k}(A)$ of a finite subset $A$ of an additive group $G$ is defined as $E_{k}(A):=|\{(a_{1},a_{2},\dots,a_{2k})\in A^{2k};\ a_{1}+a_{2}+\dots+a_{k}=a_{k+1}+a_{k+2}+\dots+a_{2k}\}|.$ The result of Kane-Tao was later extended to all $k\geqslant 2$, see [2, Theorem 2] and [12, Corollary 2], establishing that $E_{k}(A)\leqslant|A|^{\log_{2}{{2k}\choose{k}}}$ for all $A\subseteq\{0,1\}^{d}$. Combining this estimate with our Theorem 1.1 we obtain
@@
As a consequence of the simple inequality $\frac{|A+2\cdot A|}{|A|}\leqslant|A|$ and our Theorem 1.1 we obtain
@@
This can be seen as an extension of the main theorem in [3]. As in [3] the hypercube $\{0,1\}^{n}\subset\mathbb{R}^{n}$ can be projected (which is also bijection onto its image) by a transformation $T$ onto the real line $\mathbb{R}$ giving a set $P_{n}$ of cardinality $N=2^{n}$ so that any subset of $\tilde{A}\subset P_{n}$ has the same number of distinct m-sums as the set $A\subset\{0,1\}^{n}$, where $T(A)=\tilde{A}$. Thus the corollary is a direct consequence of Theorem 1.1. The projection is defined as follows: Let $n\in\mathbb{Z}^{+}$. Choose $r_{1},r_{2},\dots,r_{n}\in\mathbb{R}\setminus\mathbb{Q}$ such that $r_{i}>0$ for all $1\leqslant i\leqslant n$ and the only integer solution to $x_{1}\cdot r_{1}+x_{2}\cdot r_{2}+\cdots+x_{n}\cdot r_{n}=0,$ is $x_{1}=x_{2}=\cdots=x_{n}=0$. For every integer $0\leq j\leq n$, we define a set $P_{j}$ of $2^{j}$ real numbers, as follows. We set $P_{0}=\{1\}$. For $1\leq j\leq n$, we define $P_{j}=P_{j-1}\bigcup(r_{j}+P_{j-1})$.
@@ 2. Proof of the main results
The upper bound in Theorem 1.5 can be obtained by considering the case $A=\{0,1,\dots,k-1\}^{d}$. We will focus our attention on the lower bound. Let $\alpha:=\frac{|kA|}{|A|}$. By the Plünnecke inequality for different summands in commutative group [4, Theorem 1.1] there is a non-empty subset $B$ of $A$ such that
@@
The proof of this lemma follows from an adaptation of the proof of [5, Theorem 1.2] by Green, Matolcsi, Rusza, Shakan and Zhelezov. Similarly to [5], we obtain Lemma 2.13 by discretizing the continuous Prékopa–Leindler inequality for many functions.
@@ 3. Acknowledgments
The authors are very thankful to Jaume de Dios Pont, Pavlos Kalantzopoulos, Fedor Petrov, and Terence Tao for helpful discussions. L.B. was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813 as well as SFB 1060. P.I. acknowledges support from NSF grant CAREER-DMS-2152401.

### v2

@@ Abstract.
We show that for all $A,B\subseteq\{0,1,2\}^{d}$ we have
@@
holds for all $p\in(1,\infty)$, where $q=\frac{p}{p-1}$ is the conjugate exponent of $p$. All the estimates are dimension free with the best possible exponents. We discuss applications to various related problems.
@@ 2. Choosing $A_{1}=\ldots=A_{n}=\mathcal{C}_{m}^{d}$ we obtain
3. The main result in [21] and independently in [9] states that $p_{2,1}=\frac{\log(3)}{\log(4)}$, i.e., we have equality in (1.5) in the case $(n,m)=(2,1)$. The equality $p_{n,1}=\frac{\log(n+1)}{n\log(2)}$ was conjectured in [9], and established [^1] We are grateful to Ben Green for pointing to us Appendix B in [6] which contains more comprehensive list of references on the value $p_{n,1}$. in [2] and independently in [16]. The proof in [16] uses conformal maps and Hadamard’s three circle theorem, whereas the proof in [2] is more direct and uses a clever inductive argument. We will present a new proof (Theorem 1.4) of the equality $p_{n,1}=\frac{\log(n+1)}{n\log(2)}$ based on Descartes’ rule of signs.
@@ 4. For $m\geqslant 3$ let $\tau_{m}$ be the solution of the equation
In the next theorem we affirmatively answer the question asked in [14] which also confirms the case $m=2$ of the question in [1].
@@
$\|f\bar{*}g\|_{1}\geqslant\|f\|_{q}\|g\|_{q},\quad\text{where}\quad q=\frac{2\log(3)}{\log(5)}.$ (1.6)
@@
The inequality (1.7) follows from (1.6) by taking indicator functions of sets i.e., $f=1_{A}$ and $g=1_{B}$.
@@
The next theorem due to Landau, Logan, Shepp [16] and independently Brown, Keane, Moran and Pierce [2] states that we have equality in (1.5) for the pairs $(n,m)=(n,1)$ for all $n\geqslant 1$. We will present a new proof of it based on Decartes’ rule of signs.
Theorem 1.4 ([16], [2]).
$\|f_{1}\bar{*}\cdots\bar{*}f_{n}\|_{1}\geqslant\prod_{j=1}^{n}\|f_{j}\|_{q},\quad\text{where}\quad q=\frac{n}{\log_{2}(n+1)}.$ (1.12)
In particular, for all subsets $A_{j}\subseteq\{0,1\}^{d}$, we have
with the best possible power $p=\frac{\log_{2}(n+1)}{n}$.
A related inequality to (1.12) was established by Ivanisvili [11, Theorem 1].
@@
Our Theorem 1.5 can be seen as an extension of a result obtained by Green and Tao [8, Proposition 4.1], for $k=2$.
@@
The additive energy $E(A)$ of a finite subset $A$ of an additive group $G$ is defined as $E(A):=|\{(a_{1},a_{2},a_{3},a_{4})\in A^{4};\ a_{1}+a_{2}=a_{3}+a_{4}\}|,$ (see [20]). It was proved by Kane and Tao [13, Theorem 7] that $E(A)\leqslant|A|^{\log_{2}6}$ for all $A\subseteq\{0,1\}^{d}$. For any $k\geqslant 2$, the $k-$additive energy $E_{k}(A)$ of a finite subset $A$ of an additive group $G$ is defined as $E_{k}(A):=|\{(a_{1},a_{2},\dots,a_{2k})\in A^{2k};\ a_{1}+a_{2}+\dots+a_{k}=a_{k+1}+a_{k+2}+\dots+a_{2k}\}|.$ The result of Kane-Tao was later extended to all $k\geqslant 2$, see [3, Theorem 2] and [15, Corollary 2], establishing that $E_{k}(A)\leqslant|A|^{\log_{2}{{2k}\choose{k}}}$ for all $A\subseteq\{0,1\}^{d}$. Combining this estimate with (1.13) we obtain
@@
As a consequence of the simple inequality $\frac{|A+2\cdot A|}{|A|}\leqslant|A|$ and (1.13) one has
@@
This can be seen as an extension of the main theorem in [4]. As in [4] the hypercube $\{0,1\}^{n}\subset\mathbb{R}^{n}$ can be projected (which is also bijection onto its image) by a transformation $T$ onto the real line $\mathbb{R}$ giving a set $P_{n}$ of cardinality $N=2^{n}$ so that any subset of $\tilde{A}\subset P_{n}$ has the same number of distinct m-sums as the set $A\subset\{0,1\}^{n}$, where $T(A)=\tilde{A}$. Thus the corollary is a direct consequence of (1.13). The projection is defined as follows: Let $n\in\mathbb{Z}^{+}$. Choose $r_{1},r_{2},\dots,r_{n}\in\mathbb{R}\setminus\mathbb{Q}$ such that $r_{i}>0$ for all $1\leqslant i\leqslant n$ and the only integer solution to $x_{1}\cdot r_{1}+x_{2}\cdot r_{2}+\cdots+x_{n}\cdot r_{n}=0,$ is $x_{1}=x_{2}=\cdots=x_{n}=0$. For every integer $0\leq j\leq n$, we define a set $P_{j}$ of $2^{j}$ real numbers, as follows. We set $P_{0}=\{1\}$. For $1\leq j\leq n$, we define $P_{j}=P_{j-1}\bigcup(r_{j}+P_{j-1})$.
@@ 2. Proof of the main results
The upper bound in Theorem 1.5 can be obtained by considering the case $A=\{0,1,\dots,k-1\}^{d}$. We will focus our attention on the lower bound. Let $\alpha:=\frac{|kA|}{|A|}$. By the Plünnecke inequality for different summands in commutative group [5, Theorem 1.1] there is a non-empty subset $B$ of $A$ such that
@@
The proof of this lemma follows from an adaptation of the proof of [7, Theorem 1.2] by Green, Matolcsi, Rusza, Shakan and Zhelezov. Similarly to [7], we obtain Lemma 2.13 by discretizing the continuous Prékopa–Leindler inequality for many functions.
@@ 3. Acknowledgments
The authors are very thankful to Ben Green, Jaume de Dios Pont, Pavlos Kalantzopoulos, Fedor Petrov, and Terence Tao for helpful discussions. L.B. was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813 as well as SFB 1060. P.I. acknowledges support from NSF grant CAREER-DMS-2152401.